OpenStudy (anonymous):

what is the solution to the rational equation (5)/(x^2+3x+2)-(1)/(x+2)=(1)/(3x+3)?

OpenStudy (saifoo.khan):

factor the denominators.

OpenStudy (anonymous):

okay

OpenStudy (saifoo.khan):

\[\frac{5}{(x+1)(x+2)}-\frac{1}{x+2}=\frac{1}{3x+3}\]

OpenStudy (saifoo.khan):

Now take LCD/LCM and solve.

OpenStudy (anonymous):

okay the answer i got was 5/2 is that correct im kinda of unsure??

OpenStudy (saifoo.khan):

Correct.

OpenStudy (saifoo.khan):

I mean you're right.

OpenStudy (anonymous):

thank you so much!!!!!!